# Stochastic Optimal Control of an Industrial Power-to-Heat System with High-Temperature Heat Pump and Thermal Energy Storage

### v2

@@ Abstract
The optimal control of sustainable energy supply systems, including renewable energies and energy storage, takes a central role in the decarbonization of industrial systems. However, the use of fluctuating renewable energies leads to fluctuations in energy generation and requires a suitable control strategy for the complex systems in order to ensure energy supply. In this paper, we consider an electrified power-to-heat system which is designed to supply heat in form of superheated steam for industrial processes. The system consists of a high-temperature heat pump for heat supply, a wind turbine for power generation, a sensible thermal energy storage for storing excess heat and a steam generator for providing steam. If the system’s energy demand cannot be covered by electricity from the wind turbine, additional electricity must be purchased from the power grid. For this system, we investigate the cost-optimal operation aiming to minimize the electricity cost from the grid by a suitable system control depending on the available wind power and the amount of stored thermal energy. This is a decision making problem under uncertainties about the future prices for electricity from the grid and the future generation of wind power. The resulting stochastic optimal control problem is treated as finite-horizon Markov decision process for a multi-dimensional controlled state process. We first consider the classical backward recursion techniques for solving the associated dynamic programming equation for the value function and compute the optimal decision rule. Since that approach suffers from the curse of dimensionality we also apply Q-learning techniques that are able to provide a good approximate solution to the optimization problem within reasonable time.
Keywords:
Stochastic optimal control, Markov decision processes, Dynamic programming, Q-learning, Power-to-heat system, Renewable energy, Cost-optimal energy management
MSC:
93E20 90-08 90C40 68T05 91G60
@@ 1 Introduction
In Walden2023 , the cost-optimal operation of this electrified system with the aim of minimizing the total cost of grid power was treated as a deterministic optimization problem and solved using methods of algebraic nonconvex, nonlinear programming theory. In addition, it was assumed that future wind power generation and electricity prices were already known in advance. However, in real-world scenarios, the problem of optimal management and operation of such systems is a decision making problem under uncertainty, as precise forecasts of future wind energy supply and electricity prices are not possible. Therefore, the problem must be formulated in a stochastic framework.
In this context, a typical question needs to be addressed: At what time and at what rate should energy be stored in or withdrawn from the TES to reduce the costs of electricity drawn from the grid? To answer this question, we will treat the cost-optimal management of the underlying industrial P2H system as a stochastic optimal control problem and solve it using Markov decision process (MDP) theory.
@@
This literature is embedded in a large number of studies on optimization problems for energy systems, in particular for electrical and thermal microgrids. However, many of the articles only briefly describe the underlying model and methods. The optimization problems are mainly related to technical aspects and the control problem is solved with commercial optimization software. The mathematical aspects of the optimal control of energy systems are generally not sufficiently addressed.
@@
The optimal management of combined heating and power systems has recently been studied by several authors. In Testi et al. testi2020stochastic , an optimal integration of electrically driven heat pumps within a hybrid distributed energy supply system is investigated. There, the authors proposed a multi-objective stochastic optimization methodology to evaluate the integrated optimal sizing and operation of the energy supply systems under uncertainties in climate, space occupancy, energy loads, and fuel costs. In Kuang et al. kuang2019stochastic , a stochastic dynamic solution for off-design operation optimization of combined heating and power systems with energy storage is considered. A review on optimal energy management of combined cooling, heating and power microgrids is given in Gu et al. gu2014modeling . Further contributions on combined heating, cooling, and power system can be found in ehsan2019scenario ; zhong2021distributed and references therein.
@@
In recent years, machine learning methods have also been increasingly used to solve optimization problems. For example, Bui et al. en12091789 and Mohammed et al. DQNMicrogrid model a microgrid energy management system with battery storage that is connected to the energy grid and distributed energy sources. To obtain an optimal operation strategy that aims to handle loads, prices and the decision of charging or discharging the battery, they used Q-learning Watkins1989 . Nakabi and Toivanen NakabiTahaToivanenPekka considered a similar application, but used and compared different state-of-the-art reinforcement learning algorithms like Q-learning, deep deterministic policy gradients or proximal policy optimization to achieve the optimal management of their microgrid model. Another application of MDPs is proposed by Yu et al. DRLSmartHomeEnergyManagement , who apply reinforcement learning to a home energy management system. In addition to using a battery storage and connecting to the power grid as well as renewable energies, the household utilizes a heating, ventilation and air conditioning system that needs to be operated as cost efficient as possible. Belloni et al. Belloni2016ASO use a MDP formulation of a system with a WT and battery storage to obtain the optimal control with dynamic programming. Thermal storage devices combined with a heat pump are used in the papers of Ridder et al. and Chenzi et al. smartgreens21 , which use dynamic programming and Q-learning, respectively.
@@
The cost-optimal management of energy supply systems under uncertainty can be treated mathematically as a stochastic optimal control problem. There is extensive literature on this theory. A considerable part of this literature investigates dynamic programming solution techniques. In the continuous-time setting, in which diffusion or jump-diffusion processes form the controlled state process, this leads to the Hamilton-Jacobi-Bellman equation as a necessary optimality condition, see Fleming and Soner fleming2006control , Pham pham2009c , and Oksendal and Sulem oksendal2019stochastic . These nonlinear partial differential equations can usually only be solved using numerical methods such as those in Shardin and Wunderlich shardin2017partially , Chen and Forsyth chen2010implications . For discrete-time models, the theory of Markov decision processes provides a solution algorithm based on backward recursion. We refer here to Bäuerle and Rieder Baeuerle2011 , Puterman Puterman2014 , Hernández-Lerma and Lasserre HernandezLerma1996 and Powell Powell2011 . Such MDPs are also obtained by time discretization of continuous-time control problems.
@@
This article presents a mathematical model for the operation of an industrial P2H system with an HTHP and a TES. It explicitly takes into account the stochasticity of intermittent renewable energy sources such as wind power and the fluctuating market prices for electricity in the power grid. These variables are modeled by suitable stochastic processes that are calibrated to real data. Furthermore, the model takes into account that permanent changes in the operating points should be avoided. Therefore, the cost-optimal energy management problem is treated as a discrete-time stochastic optimal control problem, where the controls are kept constant between two discrete points in time. Nevertheless, the dynamics of the state and system variables describing the operation of the system are treated in continuous time to avoid unnecessary time discretization errors.
The optimization problem is formulated as an MDP and solved using dynamic programming methods. Since the state of the control problem is three-dimensional, the numerical solution already faces the curse of dimensionality. The problem becomes even more serious when we extend and refine our stylized model to include more details of the HTHP operation. Then the computational effort for the numerical solutions becomes prohibitively high. Therefore, in this paper, we investigate machine learning methods such as Q-learning to find a faster numerical approximate solution. Finally, we present the results of extensive numerical experiments in which we compare the results of the different numerical methods.
@@ 2.1 Industrial P2H System
The industrial P2H System based on renewable energy shown in Fig. 1 is designed to supply constant process heat in the form of superheated steam. The system consists of the following four components:
@@
The thermal system components (ii)-(iv) are connected via a thermal oil loop. A more detailed system configuration is depicted in Fig. 2. The HTHP generates high-temperature process heat, which is fed into a thermal fluid loop with thermal oil as the heat transfer fluid (HTF). Via a fluid bypass, the charging factor $l^{\text{C}}\in[0,1]$ determines the proportion of the HTF that is routed through the TES, while the remaining proportion $1-l^{\text{C}}$ passes directly into the SG. The bypass is used to regulate the charging process of the TES. During charging, the hot HTF flows through the cooler TES and heats the storage medium.
@@
Figure 2: Detailed flow diagram Walden2023 of the studied industrial P2H system (cf. Fig. 1) with HTHP, TES and SG. The charging and discharging factor $l^{\text{C}},l^{\text{D}}\in[0,1]$ determine the heat flow to the SG and the HTHP depending on the thermal state of the TES. Exemplarily, the red lines indicate the charging mode if $l^{\text{C}}\in(0,1)$, the blue lines the discharging mode for $l^{\text{D}}\in(0,1)$. Simultaneous charging and discharging is not allowed. Charging mode is characterized by $l^{\text{C}}\in(0,1],l^{\text{D}}=0$, discharging by $l^{\text{D}}\in(0,1],l^{\text{C}}=0$, and idle mode by $l^{\text{C}}=l^{\text{D}}=0$.
A second bypass from the SG outlet returning to the high-temperature heat exchanger (HTHX) of the HTHP is used to discharge the TES. The discharge factor $l^{\text{D}}\in[0,1]$ specifies the proportion of the HTF that is passed through the TES so that the remaining part $1-l^{\text{D}}$ enters directly into the HTHP. During the discharging process, the HTF cooled in the SG flows through the warmer TES and lowers the temperature of the storage medium. As the fluid now enters the HTHX at a higher temperature, the HTHP’s electricity consumption is reduced.
@@
In our setup, we use the heat flow rates that determine the charging and discharging operations of the TES to describe the control of the P2H system. These rates have a direct functional relationship with the HTHX inlet and outlet temperatures, the HTHP’s compressor shaft speed and electricity consumption, which is described in more detail below in Subsection 2.4. The HTHP electricity consumption not covered by wind energy determines the amount of electricity drawn from the grid, resulting a direct functional relationship to the running electricity costs that are included in the performance criterion of the optimization problem. More details follow below in Subsection 3.4.
The mathematical description of nonlinear component models for HTHP and SG is based on process simulation software that also takes into account the part-load behavior of the heat pump. Based on this, the physical characteristics are then approximated by algebraic surrogate models that appropriately mimic the input-output behavior of the components. For our purposes, it is sufficient to model the state of charge of the TES only by its spatially averaged temperature and neglect the detailed spatial temperature distribution, as for example in TakamWunderlichPamen2023 . This avoids complex calculation of internal heat propagation and facilitates the solution of the optimization problem.
The actual WT power output is typically modeled by a function of the wind speed at rotor height. This dependence is given by the so-called power curve, which is explained in Appendix F.3. We emphasize that the system and component modeling does not take into account pressure and heat losses as well as friction losses and also no electricity consumption of auxiliary systems such as fluid pumps. The design and dimensioning of the system components was determined by engineering calculations, for which we refer to Walden2023 .
@@ 2.2 Time Discretization
While the state and system variables of the P2H system evolve continuously over time, the control variables available to the controller are typically not changed permanently, but only at discrete points in time and then kept constant until the next time point. This is caused by the fact that operating the HTHP, i.e. changing the HTHX outlet temperature by varying the compressor shaft speed, induces thermal stresses in the heat exchangers during transient operations. For this reason, rapid changes in the operating points should be avoided and limited to a few discrete points in time.
We therefore divide the planning horizon $[0,t_{\text{E}}]$ into $N\in\mathbb{N}$ uniformly spaced subintervals of length $\Delta t={t_{\text{E}}}/{N}$ and define the time grid points $t_{n}=n\Delta t$ for $n=0,\ldots,N$. Let a continuous-time function $G:[0,t_{\text{E}}]\to\mathbb{R}$ is given, then in this work we use the short-hand notation $G_{n}:=G(t_{n})$ for the sampled value at the time grid point $t_{n}$. The control variables and some related system variables are assumed to be piecewise constant functions on the time grid introduced above and take the values $G(t)=G_{n}$ for $t\in[t_{n},t_{n+1})$ with $n=0,\ldots,N-1$. The remaining system variables and in particular the state variables of the control problem are treated as continuous-time functions governed by certain equations that capture the dynamics of the system. However, the controller only uses the values at the time grid points for the control decisions.
@@ 2.3 State and Control Variables
The exogenous states, on the other hand, are stochastic variables, as the wind speed and the electricity price are subject to a certain degree of uncertainty, meaning that future values are not known exactly in advance and are afflicted with considerable forecasting errors. They must therefore be modeled as stochastic processes, with the detailed description is being deferred to Subsection 3.1.
@@
In our model we suppose that the operation of P2H system at time $t\in[0,t_{\text{E}}]$ is controlled by
@@
with values in some action or control space $\pazocal{A}\in\mathbb{R}$ which will be specified below. In the following, we use the sign convention that a positive heat flow rate corresponds to charging, while negative values of $A$ indicate discharging. The idle mode is represented by $A(t)=0$. We show in Subsection 2.4 that specifying this variable is sufficient to adjust the other system variables describing the HTHP operation accordingly. It will also explain how the HTHP electricity demand depends on the control $A$, which in turn determines the demand for electrical energy drawn from the grid when this demand is not fully covered by wind energy, or the overproduction of wind energy that can be fed into the grid. The operational cost associated with the control $A$ are derived in Subsection 3.4. Note that the controller’s choice of the heat flow rate $A$ is subject to various constraints, which we explain in detail in Subsection 3.3.
@@
[…]
@@
$A(t)=A(t_{n})=:A_{n},\quad\text{for }~{}t\in[t_{n},t_{n+1}),~{}~{}n=0,\ldots,N-1.$ (2)
Below in Subsection 2.4, we see that the assumption of constant heat flow rates in the TES corresponds to constant oil temperatures at the inlet and outlet of the HTHX within the periods between the time grid points. This avoids too rapid changes in the HTHP operating points. Only at the time grid points $t_{n}$ the heat flow rate changes immediately from $A_{n-1}$ to $A_{n}$, whereby the transient behavior of the HTHP and TES components is neglected.
Under the above assumption that the heat flow rate $A$ is piecewise constant and does not vary within a time period $[t_{n},t_{n+1})$, the dynamics (1) of the TES temperature within such a period simplifies then to
$R(t)=R_{n}+\frac{1}{m_{S}c_{S}}A_{n}(t-t_{n}),\quad\text{and in particular }\quad R_{n+1}=R_{n}+\frac{1}{m_{S}c_{S}}A_{n}\Delta t,$ (3)

### v3

# Reinforcement Learning Methods for the Stochastic Optimal Control of an Industrial Power-to-Heat System

@@ Abstract
The optimal control of sustainable energy supply systems, including renewable energies and energy storage, takes a central role in the decarbonization of industrial systems. However, the use of fluctuating renewable energies leads to fluctuations in energy generation and requires a suitable control strategy for the complex systems in order to ensure energy supply. In this paper, we consider an electrified power-to-heat system which is designed to supply heat in form of superheated steam for industrial processes. The system consists of a high-temperature heat pump for heat supply, a wind turbine for power generation, a sensible thermal energy storage for storing excess heat and a steam generator for providing steam. If the system’s energy demand cannot be covered by electricity from the wind turbine, additional electricity must be purchased from the power grid. For this system, we investigate the cost-optimal operation aiming to minimize the electricity cost from the grid by a suitable system control depending on the available wind power and the amount of stored thermal energy. This is a decision making problem under uncertainties about the future prices for electricity from the grid and the future generation of wind power. The resulting stochastic optimal control problem is treated as finite-horizon Markov decision process for a multi-dimensional controlled state process. We first consider the classical backward recursion technique for solving the associated dynamic programming equation for the value function and compute the optimal decision rule. Since that approach suffers from the curse of dimensionality we also apply reinforcement learning techniques, namely Q-learning, that are able to provide a good approximate solution to the optimization problem within reasonable time.
@@ 1 Introduction
In Walden2023 , the cost-optimal operation of this electrified system, with the aim of minimizing the total cost of grid power was treated as a deterministic optimization problem and solved using methods of algebraic nonconvex, nonlinear programming theory. In addition, it was assumed that future wind power generation and electricity prices were already known in advance. However, in real-world scenarios, the problem of optimal management and operation of such systems is a decision making problem under uncertainty, as precise forecasts of future wind energy supply and electricity prices are not possible. Therefore, the problem must be formulated in a stochastic framework.
In this context, a typical question needs to be addressed: At what time and at what rate should energy be stored in or withdrawn from the TES to reduce the costs of grid electricity? To address this question, we will treat the cost-optimal management of the underlying industrial P2H system as a stochastic optimal control problem and solve it using Markov decision process (MDP) theory.
@@
This literature is embedded in numerous studies on optimization problems for energy systems, in particular for electrical and thermal microgrids. However, many of the articles only briefly describe the underlying model and methods. The optimization problems are mainly related to technical aspects, and the control problem is solved with commercial optimization software. The mathematical aspects of the optimal control of energy systems are generally not sufficiently addressed.
@@
The optimal management of combined heating and power systems has recently been studied by several authors. In Testi et al. testi2020stochastic , an optimal integration of electrically driven heat pumps within a hybrid distributed energy supply system is investigated. There, the authors proposed a multi-objective stochastic optimization methodology to evaluate the integrated optimal sizing and operation of the energy supply systems under uncertainties in climate, space occupancy, energy loads, and fuel costs. In Kuang et al. kuang2019stochastic , a stochastic dynamic solution for off-design operation optimization of combined heating and power systems with energy storage is considered. The cost-optimal management of a residential heating system with a geothermal storage under uncertainty is studied in Takam et al. TakamWunderlich2025 . A review on optimal energy management of combined cooling, heating and power microgrids is given in Gu et al. gu2014modeling . Further contributions on combined heating, cooling, and power system can be found in ehsan2019scenario ; zhong2021distributed and references therein.
@@
In recent years, machine learning methods have also been increasingly used to solve optimization problems. For example, Bui et al. en12091789 and Mohammed et al. DQNMicrogrid model a microgrid energy management system with battery storage that is connected to the energy grid and distributed energy sources. To obtain an optimal operation strategy that aims to handle loads, prices and the decision of charging or discharging the battery, Q-learning Watkins1989 is used. Nakabi and Toivanen NakabiTahaToivanenPekka considered a similar application, but used and compared different state-of-the-art reinforcement learning algorithms like Q-learning, deep deterministic policy gradients or proximal policy optimization to achieve the optimal management of their microgrid model. Another application of MDPs is proposed by Yu et al. DRLSmartHomeEnergyManagement , who apply reinforcement learning to a home energy management system. In addition to the usage of a battery storage and the connection to the power grid as well as renewable energies, the household utilizes a heating, ventilation and air conditioning system that needs to be operated as cost efficient as possible. Belloni et al. Belloni2016ASO use a MDP formulation of a system with a WT and battery storage to obtain the optimal control with dynamic programming. Thermal storage devices combined with a heat pump are used in the papers of Ridder et al. DERIDDER20112918 and Chenzi et al. smartgreens21 , which use dynamic programming and Q-learning, respectively.
@@
The cost-optimal management of energy supply systems under uncertainty can be treated mathematically as a stochastic optimal control problem. There is extensive literature on this theory. A considerable part of this literature investigates dynamic programming solution techniques. In the continuous-time setting, in which diffusion or jump-diffusion processes form the controlled state process, this leads to the Hamilton-Jacobi-Bellman equation as a necessary optimality condition, see Fleming and Soner fleming2006control , Pham pham2009c , and Oksendal and Sulem oksendal2019stochastic . These nonlinear partial differential equations can usually only be solved using numerical methods such as those in Shardin and Wunderlich shardin2017partially , Chen and Forsyth chen2010implications . For discrete-time models, the theory of Markov decision processes provides a solution algorithm based on backward recursion. We refer to Bäuerle and Rieder Baeuerle2011 , Puterman Puterman2014 , Hernández-Lerma and Lasserre HernandezLerma1996 and Powell Powell2011 . Such MDPs are also obtained by time discretization of continuous-time control problems.
@@
This article presents a mathematical model for the operation of an industrial P2H system with an HTHP and a TES. It explicitly takes into account the stochasticity of intermittent renewable energy sources such as wind power and the fluctuating market prices for electricity in the power grid. These variables are modeled by suitable stochastic processes that are calibrated to real-world data. Furthermore, the model takes into account that permanent changes in the operating points should be avoided. Therefore, the cost-optimal energy management problem is treated as a discrete-time stochastic optimal control problem, where the controls are kept constant between two discrete points in time. Nevertheless, the dynamics of the state and system variables describing the operation of the system are treated in continuous-time to avoid unnecessary time discretization errors.
The optimization problem is formulated as an MDP and solved using dynamic programming methods. Since the state of the control problem is three-dimensional, the numerical solution already faces the curse of dimensionality. The problem becomes even more serious when we extend and refine our stylized model to include more details of the HTHP operation. Then the computational effort for the numerical solutions becomes prohibitively high. Therefore, in this paper, we investigate reinforcement learning methods such as Q-learning to find a faster numerical approximate solution. Finally, we present the results of extensive numerical experiments in which we compare the results of the different numerical methods.
@@ 2.1 Industrial P2H System
The industrial P2H system based on renewable energy and energy storage shown in Figure 1 is designed to supply constant process heat in the form of superheated steam. The system consists of the following four components:
@@
The thermal system components (ii)-(iv) are connected via a thermal oil loop. A more detailed system configuration is depicted in Figure 2. The HTHP generates high-temperature process heat, which is fed into a thermal fluid loop with thermal oil as the heat transfer fluid (HTF). Via a fluid bypass, the charging factor $l^{\text{C}}\in[0,1]$ determines the proportion of the HTF that is routed through the TES, while the remaining proportion $1-l^{\text{C}}$ passes directly into the SG. The bypass is used to regulate the charging process of the TES. During charging, the hot HTF flows through the cooler TES and heats the storage medium.
@@
Figure 2: Detailed flow diagram Walden2023 of the studied industrial P2H system (cf. Figure 1) with HTHP, TES and SG. The charging and discharging factor $l^{\text{C}},l^{\text{D}}\in[0,1]$ determine the heat flow to the SG and the HTHP depending on the thermal state of the TES. Exemplarily, the red lines indicate the charging mode if $l^{\text{C}}\in(0,1)$, the blue lines the discharging mode for $l^{\text{D}}\in(0,1)$. Simultaneous charging and discharging is not allowed. Charging mode is characterized by $l^{\text{C}}\in(0,1],~l^{\text{D}}=0$, discharging by $l^{\text{D}}\in(0,1],~l^{\text{C}}=0$, and idle mode by $l^{\text{C}}=l^{\text{D}}=0$.
A second bypass from the SG outlet returning to the high-temperature heat exchanger (HTHX) of the HTHP is used to discharge the TES. The discharge factor $l^{\text{D}}\in[0,1]$ specifies the proportion of the HTF that is passed through the TES, such that the remaining part $1-l^{\text{D}}$ enters directly into the HTHP. During the discharging process, the HTF cooled in the SG flows through the warmer TES and lowers the temperature of the storage medium. As the fluid now enters the HTHX at a higher temperature, the HTHP’s electricity consumption is reduced.
@@
In our setup, we use the heat flow rates that determine the charging and discharging operations of the TES to describe the control of the P2H system. These rates have a direct functional relationship with the HTHX inlet and outlet temperatures, the HTHP’s compressor shaft speed and electricity consumption, which is described in more detail below in Subsection 2.4. The HTHP electricity consumption that is not covered by wind energy determines the amount of electricity drawn from the grid, resulting in a direct functional relationship with the running electricity costs that are included in the performance criterion of the optimization problem. More details follow below in Subsection 3.4.
The mathematical description of nonlinear component models for HTHP and SG is based on process simulation software that also takes into account the part-load behavior of the heat pump. Based on this, the physical characteristics are then approximated by algebraic surrogate models that appropriately mimic the input-output behavior of the components. For our purposes, it is sufficient to model the state of charge of the TES only by its spatially averaged temperature and neglect the detailed spatial temperature distribution, as for example in TakamWunderlichPamen2023 ; TakamWunderlich2025Energies . This avoids complex calculation of internal heat propagation and facilitates the solution of the optimization problem.
The actual WT power output is typically modeled by a function of the wind speed at rotor height. This dependence is given by the so-called power curve, which is explained in Subsection Appendix G.3. We emphasize that the system and component modeling does not take into account pressure and heat losses as well as friction losses and also no electricity consumption of auxiliary systems such as fluid pumps. The design and dimensioning of the system components was determined by engineering calculations, for which we refer to Walden2023 .
@@ 2.2 Time Discretization
While the state and system variables of the P2H system evolve continuously over time, the control variables available to the controller are typically not changed permanently, but only at discrete points in time and then kept constant until the next time point. This is caused by the fact that operating the HTHP, i.e., changing the HTHX outlet temperature by varying the compressor shaft speed, induces thermal stresses in the heat exchangers during transient operations. For this reason, rapid changes in the operating points should be avoided and limited to a few discrete points in time.
We therefore divide the planning horizon $0,t_{\text{E}}]$ into $N\in\mathbb{N}$ uniformly spaced subintervals of length $\Delta t={t_{\text{E}}}/{N}$ and define the time grid points $t_{n}=n\Delta t$ for $n=0,\ldots,N$. Let $G:[0,t_{\text{E}}]\to\mathbb{R}$ be a given continuous-time function. In this work we use the short-hand notation $G_{n}:=G(t_{n})$ for the sampled value at the time grid point $t_{n}$. The control variables and some related system variables are assumed to be piecewise constant functions on the time grid introduced above and take the values $G(t)=G_{n}$ for $t\in[t_{n},t_{n+1})$ with $n=0,\ldots,N-1$. The remaining system variables and in particular the state variables of the control problem are treated as continuous-time functions governed by certain equations that capture the dynamics of the system. However, the controller only uses the values at the time grid points for the control decisions.
@@ 2.3 State and Control Variables
The exogenous states, on the other hand, are stochastic variables, as the wind speed and the electricity price are subject to a certain degree of uncertainty, meaning that future values are not known exactly in advance and are afflicted with considerable forecasting errors. They must therefore be modeled as stochastic processes, where the detailed description is deferred to Subsection 3.1.
@@
In our model, we suppose that the operation of P2H system at time $t\in[0,t_{\text{E}}]$ is controlled by
@@
with values in some action or control space $\pazocal{A}\in\mathbb{R}$ which will be specified below. In the following, we use the sign convention that a positive heat flow rate corresponds to charging, while negative values of $A$ indicate discharging. The idle mode is represented by $A(t)=0$. We show in Subsection 2.4 that specifying this variable is sufficient to adjust the other system variables describing the HTHP operation accordingly. There it will also be explained how the HTHP electricity demand depends on the control $A$, which in turn determines the demand for electrical energy drawn from the grid, when this demand is not fully covered by wind energy, or the overproduction of wind energy that can be fed into the grid. The operational cost associated with the control $A$ are derived in Subsection 3.4. Note that the controller’s choice of the heat flow rate $A$ is subject to various constraints, which we explain in detail in Subsection 3.3.
@@
The control $A$ directly determines the dynamics of the only controlled/endogenous state variable in our control system, namely the TES temperature $R$. It results from an energy balance that describes the change in thermal energy in the TES due to the inflow and outflow of energy during charging and discharging. Note, that we neglect thermal losses to the environment. Then the change in thermal energy in the time interval $t_{a},t_{b}]$ with $0\leq t_{a}<t_{b}\leq t_{\text{E}}$ in the continuous-time setting is given by $\int_{t_{a}}^{t_{b}}A(s)\,\mathrm{d}s$. On the other hand, it is equal to $m_{\text{s}}c_{\text{p,s}}(R(t_{b})-R(t_{a}))$, where $c_{\text{p,s}}$ and $m_{\text{s}}$ denote the specific heat capacity (assumed to be temperature independent) and the mass of the storage medium, respectively. This leads to the following relation for the TES temperature for $t\in[t_{a},t_{b}]$
$R(t)=R(t_{a})+\frac{1}{m_{\text{s}}c_{\text{p,s}}}\int_{t_{a}}^{t_{b}}A(s)\,\mathrm{d}s.$ (1)
@@
$A(t)=A(t_{n})=:A_{n},\quad t\in[t_{n},t_{n+1}),~n=0,\ldots,N-1.$ (2)
Below in Subsection 2.4, we notice that the assumption of constant heat flow rates in the TES corresponds to constant oil temperatures at the inlet and outlet of the HTHX within the periods between the time grid points. This avoids rapid changes in the HTHP operation. Only at the time grid points $t_{n}$ the heat flow rate changes immediately from $A_{n-1}$ to $A_{n}$, whereby the transient behavior of the HTHP and TES components is neglected.
Under the above assumption that the heat flow rate $A$ is piecewise constant and does not vary within a time period $t_{n},t_{n+1})$, the dynamics (1) of the TES temperature within such a period simplifies then to
$R(t)=R_{n}+\frac{1}{m_{\text{s}}c_{\text{p,s}}}A_{n}(t-t_{n}),\quad\text{and in particular }\quad R_{n+1}=R_{n}+\frac{1}{m_{\text{s}}c_{\text{p,s}}}A_{n}\Delta t,$ (3)